Find the directional derivitive of f(x, y, z) xy at the point (4,1,1) that goes in V=2i-j+2k direction.
Answer #1

D f (4,1,1) = f (4,1,1). zł 29 ze = z î - 9 + 28 120) = x 4+1+4 3 ng Z2 of = (for, fo, fz) = ( 3 2 2 2 Of (4,),1) = (1,4,- 4)